OpenStudy (anonymous):

Write the equation of a line that passes through (0, 4) and has a slope of -1/5

OpenStudy (akashdeepdeb):

Do you know this form of a line? y = mx + c [Where m is the slope and c is the y-intercept] Use that here. :)

OpenStudy (akashdeepdeb):

WELCOME TO OPENSTUDY! :D

OpenStudy (issy14):

You can either utilize the slope intercept form described by akashadeepDeb or the point slope formula : y-y1= m(x-x1). Where you will input your ordered pairs in y1 , m=slope, and x1.

OpenStudy (issy14):

Thanks